### v1

# Stick to your Role! Stability of Personal Values Expressed in Large Language Models

Abstract
The standard way to study Large Language Models (LLMs) through benchmarks or psychology questionnaires is to provide many different queries from similar minimal contexts (e.g. multiple choice questions). However, due to LLM’s highly context-dependent nature, conclusions from such minimal-context evaluations may be little informative about the model’s behavior in deployment (where it will be exposed to many new contexts). We argue that context-dependence should be studied as another dimension of LLM comparison alongside others such as cognitive abilities, knowledge, or model size. In this paper, we present a case-study about the stability of value expression over different contexts (simulated conversations on different topics), and as measured using a standard psychology questionnaire (PVQ) and a behavioral downstream task. We consider 19 open-sourced LLMs from five families. Reusing methods from psychology, we study Rank-order stability on the population (interpersonal) level, and Ipsative stability on the individual (intrapersonal) level. We explore two settings: with and without instructing LLMs to simulate particular personalities. We observe similar trends in the stability of models and model families - Mixtral, Mistral and Qwen families being more stable than LLaMa-2 and Phi - over those two settings, two different simulated populations, and even in the downstream behavioral task. When instructed to simulate particular personas, LLMs exhibit low Rank-Order stability, and this stability further diminishes with conversation length. This highlights the need for future research directions on LLMs that can coherently simulate a diversity of personas, as well as how context-dependence can be studied in more thorough and efficient ways. This paper provides a foundational step in that direction, and, to our knowledge, it is the first study of value stability in LLMs. The project website with code is available at https://sites.google.com/view/llmvaluestability.
Keywords: Large Language Models; LLM Stability; Context-dependence; Personal Values; Simulated populations
1 Introduction
In recent years, there has been an emergence of research using psychological tools to study Large Language Models (LLMs). In those studies, LLMs have often been used to simulate populations by instructing them to simulate different personas (?, ?). LLMs have also been used without providing such instructions, i.e. treated as a participant in a human study (?, ?). Questions in such studies have revolved around how Language Models express values (?, ?), personality traits (?, ?, ?), cognitive abilities (?, ?), and how they could replicate human data (?, ?).
Figure 1: How do LLM’s expressed values change as a function of context? An LLM is first prompted to play a specific role (e.g. Gandalf). Then, a conversation on a topic (e.g. joke) with an interlocutor model (same LLM prompted to simulate a human user) is generated. Then, the LLM simulating the persona is given a psychology questionnaire aimed to assess its expressed values. We study the stability of these expressed values across diverse conversation topics and lengths. We consider various personas to be simulated, as well as the case when the LLM is not prompted to play any particular persona. The messages and instructions in gray are set manually, and the messages in white are generated.
The use of psychological tools with LLMs opens up many scientific questions, for example regarding the nature of how the text generated by LLMs depends on context, i.e. information present in prompts or prior interaction with the model. Prompts or prior interaction, which we denote here as ’context’, can include any textual information such as instructions (e.g. personas to simulate), the dialogue history, stories written in specific styles, etc. Such contexts guide the generation of text by LLMs: different contexts may result in the expression of different behavior and values. This is sometimes beneficial and expected, e.g. an instruction to simulate some persona should influence the behavior and expressed values to be more aligned with that persona. However, this can also be problematic, e.g. a specific conversation topic may drastically influence the expressed behavior and values in unexpected ways as we will highlight in experiments below. It should be noted that, depending on the application, different types of context-dependence can be beneficial or not.
The issue of unexpected context dependence in LLM is of crucial importance. Standard evaluation benchmarks, including those using psychological questionnaires to assess properties of LLMs, consist of sets of queries, often presented with a similar minimal context (e.g. knowledge or value-related questions presented as multiple choice questions with limited context). When deployed, LLMs are exposed to many new unforeseen contexts. This means that the standard benchmarks, by themselves, cannot estimate a model’s behavior in deployment (due to the LLMs’ highly context dependent nature). It is therefore crucial to evaluate the robustness of different models to unexpected context-based changes in behavior.
This challenge is particularly acute with the use of psychological questionnaires aimed at measuring psychological dimensions like values. Those tools were initially designed to probe humans, and thus make various assumptions about humans: for example, it is expected that the answers of most humans to questionnaires about value preferences should not be significantly influenced by the content of a randomly picked Wikipedia article shown to them beforehand. As we will show below, such an assumption does not hold for many LLMs, and thus strongly limits general interpretability of using these questionnaires in a context-independent manner. It is thus key to understand better how LLMs’ behavior (e.g. expression of values) may maintain coherence or change as a function of various kinds of contexts (ranging from explicit instruction to play a particular persona to discussions about topics that seem unrelated to the expressed psychological dimensions one studies).
Previous research included certain experiments regarding unwanted context-based change (usually regarding syntactic changes in the prompt). These experiments led to conflicting results, sometimes showing robustness (?, ?, ?, ?, ?), and sometimes sensitivity (?, ?, ?). These inconclusive results motivate research into the nature and the extent of context-dependence of various LLMs.
In this paper, we present a case-study focusing on studying the stability of value expressed in 19 open-source LLMs from 5 families. First, we study to what extent can LLMs simulate various personas in coherent ways, i.e. expressed values should change according to the instructed persona, but not based on the topic of a conversation not related to values. In other words, we study to what extent do LLMs’ value profiles change in wanted ways (i.e. based on an instruction), while remaining robust to unwanted context-based change (i.e. based on different conversation topics). We instruct LLMs to simulate two populations: fictional characters and well-known real-world personas. In addition, we also study the coherence and robustness of LLMs’ value expression when they are not instructed to simulate any specific persona, corresponding to frequent real-world use cases. To our knowledge, this is the first study on value stability in LLMs.
We use the Schwartz’s theory of Basic Personal Values (?, ?) and the corresponding Portrait Values Questionnaire (PVQ-40) (?, ?). This theory and questionnaire has been thoroughly studied and validated in the field of psychology, and it outlines ten basic personal values (Universalism, Benevolence, Conformity, Tradition, Security, Power, Achievement, Hedonism, Stimulation, and Self-Direction). Following that research, we outline two types of value stability studies in the psychology literature: Rank-Order (on a population/interpersonal level) and Ipsative (on an individual/intrapersonal level), see details below and on figures 2 and 3.
The main contributions of this paper are:
  • Introduction and adaptation of the methodology for evaluating Rank-Order and Ipsative stability in LLMs
  • First analysis of value stability across contexts (including various conversation lengths)
  • Systematic comparison of Rank-Order and Ipsative stability of 19 LLMs with and without instructing the models to simulate specific personas
  • Analysis of stability in a downstream behavioral task
Figure 2: Rank-Order stability An example of estimating Rank-Order stability for benevolence. In each context, characters are ordered according to their benevolence scores in that context. In this example, the orders are almost the same in contexts 1 and 2 (high Rank-Order stability), and very different in contexts 2 and 3 (low Rank-Order stability).
Figure 3: Ipsative stability An example of estimating Ipsative stability for a character (Gandalf). Values are ordered according to the character’s scores in each context. In this example, the orders are the same in contexts 1 and 2 (high Ipsative stability), and different in contexts 2 and 3 (low Ipsative stability).
2 Methods
This section discusses how we administer the PVQ questionnaire over different contexts to evaluate value stability. We conduct experiments in two ways: with and without instructing the models to simulate specific personas. Different contexts are induced by simulating conversations on different topics with a separate instance of the same model (the interlocutor). To set a conversation topic (e.g. joke) we manually set the first interlocutor’s message (e.g. "Tell me a joke."). We let the models exchange $n$ messages, manually set the last interlocutor’s message as the query (e.g. PVQ item), and record the model’s response. After the questionnaire was given in each context, we estimate the Ipsative and Rank-Order stability. This process is shown in figure 1.
2.1 Administering the questionnaire
Administering the questionnaire consists of the following steps depicted in figure 1:
1. A model is instructed to simulate a persona (optional). We study personas from two populations: 1) 60 fictional characters from J.R.R. Tolkien’s universe , and 2) 50 real world personas (see appendix A for details). A persona setting instruction (e.g. “You are Gandalf from J. R. R. Tolkien’s Middle-earth legendarium.”) is given to the model (see appendix A for details).
2. A separate interlocutor model instance is created. The interlocutor model is an instance of the same model as the one being tested. The interlocutor is given the following instruction: “You are simulating a human using a chatbot. Your every reply must be in one sentence only.” If a persona was provided in step 1, the following sentence is added as the second sentence: “The chatbot is pretending to be character_name.”
3. A conversation topic is induced. The first interlocutor’s message is manually set to induce one of the following topics: grammar, joke, poem, history, chess. For example, to induce the topic of “joke” it is set to “Tell me a joke.”. See appendix A for details.
4. A conversation is simulated. The two models are let to exchange $n$ messages. In our experiments, $n$ is set to 3 (except when studying the effect of $n$ on stability).
5. A questionnaire is given. A questionnaire item is manually set as the interlocutor’s last message with a random order of suggested answers, and the model’s response is recorded. This is repeated for each question in parallel (with the same simulated conversation). That way, the model’s response is not influenced by responses to other questions.
6. A questionnaire is scored. The responses are scored to obtain the scores for the ten values. See appendix A for details.
7. Stability is estimated. If a persona was provided in step 1, steps 1 to 6 are repeated for every persona in the simulated population. Then, the whole process is repeated with five random seeds. Stability is estimated for each seed and then averaged, i.e. value stability for one model is estimated from 50/60k queries, depending on the population (5 seeds x 5 topics x 50/60 personas x 40 PVQ items). For reference, MMLU (?, ?) (a commonly used general knowledge benchmark) contains 14k test questions.
If no persona was provided in step 1, steps 2 to 6 are repeated 50 times with different seeds. As no persona was provided, this process repeats the same experiment with 50 different permutations in the order of suggested answers, and therefore no additional seeds are needed. Ipsative stability is computed for each of the 50 permutations and then averaged, i.e. value stability for one model is estimated from 10-12k queries (5 topics x 50/60 permutations x 40 PVQ items).
2.2 Estimating the stability
We estimate two types of value stability: Rank-Order and Ipsative. Rank-Order estimates the stability of some value at the population (interpersonal) level, as the stability of the order of participants in expressing that value. Intuitively, this can be seen as addressing the following question: "Does Jack always value Tradition more than Jane does?". Ipsative stability estimates the stability at the individual (intrapersonal) level as the stability of individuals value hierarchies. Intuitively, this can be seen as addressing the following question: "Does Jack always value Tradition more than Benevolence?".
Rank-Order stability
Rank-Order stability is used to estimate the stability of some value inside a population. In psychology, Rank-Order stability for some value can be computed as the correlation in the order of individuals’ scores at two points in time (Spearman correlation between the participants’ ranks). Here, instead of comparing the participant ranks at two points in time, we compare it in different contexts (see Fig. 2). We evaluate a model in five different contexts and compute the stability for each pair of contexts, and estimate the final stability as the average of those pairs.
Ipsative (within-person) stability
Ipsative stability is used to estimate the stability of an individual’s value profile. In psychology, Ipsative stability can be computed as the correlation between the ranks of values for the same individual at two points in time (Spearman correlation between the values’ scores). Here, instead of evaluating the value profile at two points in time, we evaluate it in different contexts (see Fig 3).
We evaluate models in five different contexts and compute the stability between each pair of contexts. We estimate the final stability by averaging over those pairs.
3 Experiments
This section provides an analysis of Ipsative and Rank-Order stability in LLMs. LLMs will be evaluated in two ways: with and without instructing the models to simulate particular personas. We aim to address the following questions:
  • How do different models and model families compare in terms of expressed value stability?
  • How does the stability of values expressed by LLMs compare to stability observed in human development?
  • Can LLMs keep coherent personas over longer conversations?
  • To what extent do conclusions made with PVQ transfer to a downstream behavioral task?
3.1 Models
The LLaMa-2 (?, ?) family contains models with 7, 13 and 70 billion parameters. It also includes “chat” versions, which were fine-tuned on instructions and with RLHF (?, ?). The Mistral (?, ?) family contains base and instruction fine-tuned models with 7 billion parameters. Zephyr (?, ?) also belongs in this family as a DPO (?, ?) tuned version of the base Mistral model. The Mixtral (?, ?) family contains base and “instruct” models with 46.7 billion parameters. The “instruct” version was trained by supervised fine-tuning and DPO (?, ?). We consider these two models and their 4-bit quantized versions. The Phi (?, ?) family contains smaller base models, of which we consider two models with 1.3 and 2.7 billion parameters. From the Qwen (?, ?) model family we consider base models with 7, 14, and 74 billion parameters.
((a)) Rank-order value stability for LLMs simulating fictional characters.
((b)) Rank-order value stability for LLMs simulating real-world personas.
((c)) Rank-order donation stability on a downstream behavioral task for LLMs simulating fictional characters.
Figure 4: Rank-order stability ($Mean\pm SI(\alpha=0.05)$) of personal values and donations exhibited by LLMs following conversations on different topics. Consistent trends are visible: Mixtral, Qwen, and Mistral model families are the most stable, compared to LLaMa-2 and Phi families. All models exhibit lower than human stability, despite the comparison being skewed in their favor.
Figure 5: Ipsative value stability ($Mean\pm SI(\alpha=0.05)$) of LLMs without the persona setting instructions. Mistral-Instruct-v0.1 and Qwen-72B models show the highest stability. Mixtral, Mistral and Qwen families are more stable. Human change is shown for reference, but no strong conclusions can be made because the comparison is skewed in the LLMs’ favor.
3.2 How do different models and model families compare in terms of expressed value stability?
We evaluate the Rank-Order stability of LLMs instructed to simulate various personas. Figure 4 compares models’ value stability of two simulated populations: fictional characters (Fig. 3(a)) and real-world personas (Fig. 3(b)). On fictional characters the most stable model is Mixtral-Instruct-v1 ($r=0.42$), which is followed by its 4-bit quantized version ($r=0.32$), Mistral-Instruct-v2 ($r=0.31$), and Qwen-72B ($r=0.26$). A similar trend is observed on real-world personas, however, Qwen-72B ($r=0.5$) matches the stability of Mixtral-Instruct-v1 ($r=0.51$). More generally, we observe consistent trends in terms of model families in both simulated populations: Mixtral, Mistral and Qwen families show more stability than LLaMa-2 and Phi.
Figure 5 compares the Ipsative stability of LLMs without instructing them to simulate any particular persona. While similar trends of models are observed to those in the Rank-Order experiments, the models are less polarized. While Mixtral-Instruct-v0.1 ($r=0.84$), its 4-bit quantized version ($r=0.8$), and Qwen-72B ($r=0.78$) are again the most stable models, zephyr-7b-beta ($r=0.62$) is more stable than Mistral-Instruct-v2 ($r=0.45$). Furthermore, compared to the previous experiment, stability is also observed in LLaMa-2-70b-chat ($r=0.47$) and to a lesser extent in Phi-2 ($r=0.29$), LLaMa-2-70b ($r=0.17$), and Qwen-7B ($r=0.22$). The most stable model families are again Mixtral, Mistral and Qwen.
Instruction or chat fine-tuning seems to be beneficial for Ipsative stability, as every tuned model in Fig. 5 is more stable than its base version. This effect is not as conclusive for Rank-order stability. As fine-tuning adapts the model towards instruction following, dialogues (chat textual format), and answering questions, it is expected to increase stability. However, it also often includes “aligning” the model by making it less prone to exhibit unwanted behavior, which can have a detrimental effect on simulating some personas such as villains. We hypothesize that this is the reason why we observe a consistent effect only on Ipsative stability.
3.3 How does the stability of values expressed by LLMs compare to stability observed in human development?
To get a more intuitive impression of the observed stability levels, we extract data from two longitudinal studies on humans. ? (?) followed 20-year-olds for eight years and ? (?) followed 10-year-old for 2 years (these changes are denoted by horizontal lines in figures 4 and 5). It is important to note that this comparison is skewed in the LLMs favor. It is easier for LLMs to show stability in the following ways: 1) human value changes were caused by much more drastic circumstances (years of development compared to topic change in LLMs) 2) the human population was more unstable (10-year-old and 20-year-olds compared to well-established fictional characters or real-world personas). Therefore, an argument can only be made in one direction: if some models show lower stability than that observed in humans, those models can be said to exhibit subhuman value stability.
Figures 3(a) and 3(b) show that all models, when instructed to simulate various personas, exhibit much lower Rank-order stability than that observed in human populations ($r=0.57$ for ages 10 to 12, and $r=0.66$ for ages 20 to 28). The fact that LLMs show lower stability despite the comparison being skewed in their favor shows that LLMs exhibit sub-human value stability and are significantly more susceptible to unexpected context changes. These results motivate reserach on LLMs focused on simulating populations.
Figure 5 shows the Ipsative stability of models that were not instructed to simulate a persona. Both Mixtral-Instruct models ($r=0.84$ and $r=0.8$), Qwen-72B ($r=0.78$), and zephyr-7b-beta ($r=0.62$) do not exhibit lower stability than that observed in humans ($r=0.66$ for ages 10 to 12, and $r=0.59$ for ages 20 to 28)). Crucially, as discussed above, this does not imply that those models show human-level value stability, rather, the only insight is that other models show very low Ipsative stability.
Figure 6: Rank-order value stability ($Mean\pm SI(\alpha=0.05)$) of the Mixtral-Instruct model simulating fictional characters. Stability decreases with longer simulated conversations.
Figure 7: Ipsative value stability ($Mean\pm SI(\alpha=0.05)$) of LLMs with (Mixtral-Instruct) and without persona setting instructions. All models retain the same stability level in longer conversations.
3.4 Can LLMs keep coherent value profiles over longer conversations?
In the previous experiment, models were let to exchange $n=3$ messages (not including the manually set first and last interlocutor’s messages). Here, we evaluate the effect of simulated conversation length ($n$) on value stability.
Figure 6 shows the effect of simulated conversation length on Rank-order stability expressed by the Mixtral-Instruct model instructed to simulate fictional characters. Due to computational constraints (evaluating a model on one population requires 300k queries), we conduct this experiment only with Mixtral-Instruct (the most stable model from Fig. 3(a)), and only on one population. We can see that, even for this most stable model, stability diminishes as conversations get longer. This experiment highlights the limitations of LLMs in maintaining coherent interpersonal value profiles over longer conversations.
Figure 7 shows the effect of conversation length on Ipsative stability. We compare the most stable models from Fig. 5 without persona instructions, and Mixtral-Instruct with instructions to simulate fictional characters. Ipsative stability remains stable regardless of the simulated conversation length for all models. Mixtral-Instruct with persona instruction ("Mixtral-Instruct (fict. char.)"), while highly stable, is less stable than uninstructed Mixtral-Instruct ("Mixtral-Intruct"). This implies that Mixtral-Instruct is slightly better adapted for use without the persona instructions.
Mixtral-Instruct with persona instructions exhibits a combination of low Rank-Order stability (Fig. 6) and high Ipsative stability (Fig. 7). This implies that the model is able to maintain characters consistently, however those characters might not be adequately impersonated (e.g. simulating different characters with similar value profiles). This hypothesis is confirmed in Appendix B.2. These results suggest that current LLMs are not well suited for use with persona setting instructions, and motivate future research on LLMs focused on simulating specific personas.
3.5 To what extent do conclusions made with PVQ transfer to a downstream behavioral task?
In this experiment, we study if the conclusions made with the PVQ questionnaire transfer to a downstream behavioral task (if models that exhibited more stable value profiles also exhibit more stable behavior on a downstream task). We construct a task where an LLMs can choose an amount of coins (0 to 10) to give a beggar. The full test set consists of 100 queries with beggars of different names, genders, and fictional races (elves, dwarves, orcs, humans, hobbits). The average amount of donated coins is computed for each race. The stability of donated coins is then estimated in the same way as value stability, i.e. amounts donated to different races are treated in the same way as scores for different values. (see appendix A.6).
Figure 3(c) compares models according to their Rank-order stability on the donation task. The base Mixtral model (Mixtral-Base) shows the highest stability ($r=0.32$), and is closely followed by Mixtral-Instruct ($r=0.26$). Those models are then followed by Qwen-72B ($r=0.23$), the 4-bit quantized Mixtral base ($r=0.22$) and instruct ($r=0.18$) models, and then by Mistral-Instruct-v0.2 ($r=0.20$).
Compared to the results on PVQ (Fig. 3(a)), the overall trend of models is consistent with some differences. Qwen-72B, Mixtral-Instruct, and Mistral-Instruct-v0.2 are again among the most stable models. However, the Mixtral base models not only show stability, but even outperform their Mixtral-Instruct counterparts. Furthermore, zephyr-7b-beta shows lower, and the LLaMa-2-70b base model shows higher stability. The trends of model families are consistent: Mixtral, Mistral and Qwen are again the most stable, while Phi and LLaMa-2 show low stability. This experiment shows that, while the overall general trends are consistent between PVQ and the donation task, there are some differences (most notably the high stability of the base Mixtral model).
In appendix B.3 we study if behavior correlates with value expression, i.e. if donation amounts correlate positively with the expression of benevolence and universalism, and negatively with the expression of power and achievement.
4 Conclusion
This paper presents the first study into the stability of values expressed by Large Language Models. We consider (interpersonal) Rank-Order stability and (intrapersonal) Ipsative stability. We evaluate value stability over different contexts induced by simulating conversations about different topics. We conduct experiments with and without instructing the models to simulate particular personas. Over our experiments, we observed consistent trends of value stability: Mixtral, Mistral and Qwen model families were more stable (these trends are also confirmed on a downstream behavioral task). LLMs instructed to simulate personas exhibit much lower than human stability (despite the comparison being skewed in their favor), which further diminishes over longer conversations. This insight motivates future research on models specialized in simulating coherent populations of individuals.
This paper highlights how seemingly unrelated context changes can result in unpredictable and unwanted changes in behavior. We argue that context-dependence, and more precisely, value stability, should be seen as another dimension of LLM comparison alongside knowledge, model size, speed, and similar. Instead of evaluating LLMs with many different questions from a single minimal context, they should also be evaluated (in terms of their context-dependence and value stability) with the same questions asked in many different contexts. This study presents a first step in that direction.
Limitations Due to computational requirements for evaluating LLMs, the evaluation of stability considers only five different conversation topics and simulated rather short conversations (up to 9 messages). Greatly increasing the number of topics and conversation length could provide more precise insights into the stability of various models. For one model, Rank-Order value stability is estimated from 50/60k queries and Ipsative stability from 10/12k queries (depending on the simulated population). For reference, MMLU (?, ?) (a commonly used general knowledge benchmark) contains 14k test questions.
Future work We believe that this paper opens many research avenues regarding context-dependence and value stability of LLMs. Similar questions to those explored in this paper could be explored for personality traits, cultural values, cognitive abilities and knowledge. An interesting direction is to explore if the same model can exhibit high stability in both settings with and without the persona instruction, or if specialized models are required. This paper opens a new area of research in creating, evaluating and analyzing models specialized in simulating coherent and diverse populations. Such models are needed for many applications such as replicating human studies (?, ?), simulating social interactions (?, ?), training teachers (?, ?), and many more.
5 Acknowledgments
Experiments presented in this paper were carried out using the HPC resources of IDRIS under the allocation 2023-[A0151011996] made by GENCI. We would also like to thank Cédric Colas for many helpful discussions.
Appendix A Additional details on the methods
A.1 Constructing the simulated populations
We construct two simulated populations: Fictional characters and Real-world personas. Instructing a model to simulate a fictional or real persona that is well-known is a simple way to precisely define a persona’s character (assuming the models were trained on enough data related data). A list of personas from both populations is shown in table 1.
The Fictional characters population consists of characters from J.R.R. Tolkien’s universe. They were initially selected based on the length of their Wikipedia page (?, ?). Then, some characters were manually replaced by including more female characters and villains to make the population more balanced. Here is an example of an instruction inducing a persona from this population: “You are Gandalf from J. R. R. Tolkien’s Middle-earth legendarium.” This population contains a total of 60 fictional characters, which are listed in table 1.
The Real-world personas population contains personas from an online list of influential people (?, ?). Here is an example of an instruction inducing a persona from this population: “You are Marilyn Monroe (1926 – 1962) American actress, singer, model.” This population contains a total of 50 personas, which are listed in table 1.
A.2 Formatting the prompt
The full prompting procedure consists of five parts depicted by different colors in figure 11. The model is instructed to simulate a persona (blue), a conversation topic is induced (green), and a conversation is simulated (orange). The conversation is simulated using a separate instance of the same model: The Interlocutor model. This instance is given the instruction to simulate a human user (see section 2 and Fig. 1). Then, a query is posed (red), followed by a query string. The query string skews the next token distribution towards capital letters denoting an answer. The model generates a distribution for the next token, and the answer is taken as the most probable token from a set of relevant capital letters (e.g. A to F).
The precise prompt format depends on the LLM. We consider three types: 1) Base models: LLaMa-[7|13|70]b, Mi[s|x]tral-Base, phi-[1|2], Qwen-[7|14|72]B; 2) Tuned models with a system message input: LLaMa-[7|13|70]b-chat, zephyr-7b-beta; 2) Tuned models without a system message input: Mistral-Instruct-v0.[1|2], Mi[x]stral-Instruct. This specific system input is intended to provide more salient instructions to the model.
The base model format is depicted in figure 10(a). The persona is inducted to the Tested model with the following phrase: “CONTEXT: The following is a conversation with <persona>” (blue). Furthermore, the messages from the tested model are prefixed by “<persona_name>:” and from the interlocutor are prefixed by "USER:" (orange and purple). Hence, the tested model is modeling a conversation between a user and the persona. The Interlocutor model is prompted using the same format. It is given the following instruction: “CONTEXT: The following is a conversation between a human and a chatbot. The chatbot is pretending to be <persona_name>. The human’s every reply must be in one sentence only.”. The interlocutor model’s messages are prefixed by "HUMAN:", and the tested model’s responses by "<persona_name> (CHATBOT)". Hence, the interlocutor model is modeling a conversation between a human user and an LLM simulating a persona.
The formats for model tuned with and without the system message input are depicted in figure 10(b) and 10(c) respectively. Each tuned model has a specific template defining which parts of the text are messages from the user and which parts were generated by the model. For example, the Mistral model defines the user instruction between “[INST]” and “[/INST]” tags. In figures 10(b) and 10(c), keywords <system>, <user> and <assistant> denote the model specific tags. The only difference in format for the two types of models is the persona induction (blue). For models that have a system input, the persona inducing instruction (i.e. “You are <persona>”) is given through the system input. For models without such an input, the instruction is given as a user message followed by a manually set acknowledgment. 10(c) in blue). For both tuned model types, the same prompt format is used for the interlocutor model with the following instruction: "You are simulating a human using a chatbot. The chatbot is pretending to be <persona_name>. Your every reply must be in one sentence only."
A.3 Inducing a conversation topic
We induce a conversation topic by manually setting the first message of the interlocutor model. We consider the following five topics and their corresponding messages: 1) poem: “Hello, let’s write a poem together. You start by the first verse I’ll add the second one, and so on.”, 2) joke: “Tell me a joke.”, 3) history: “What is the significance of the battle of Hastings. Answer in two sentences.", 4) chess: "1. e4”, 5) grammar: “Can you check this sentence for grammar? Whilst Jane was waiting to meet hers friend their nose started bleeding.”.
A.4 Querying the model and evaluating its response
Following the simulated conversation, the model is presented with a query from a questionnaire or from the downstream donation task. The PVQ questionnaire is available in versions for male and female participants. We give the PVQ version according to the simulated persona’s gender. In experiments where no persona is simulated, and the questionnaire is given 50 times with different permutations in the order of suggested answer, male and female versions are used for 25 permutations each.
Following the full prompting procedure explained above and depicted in figure 11, a model generates a distribution for the next token. We score the probability of tokens corresponding to suggested answers. For the 6 point Likert scale, this corresponds to capital letters from A to F. The most probable letter is taken as the model’s answer.
A.5 Scoring the PVQ questionnaire
Once the answers have been collected, the scoring proceed as described in the questionnaires manual. The answers are coded from 1 to 5, where “Not like me at all” corresponds to 1 and “Very much like me” to 5. The per-participant mean score is subtracted from those scores to normalize the answers around 0. This step is crucial, as personal values are defined in relative terms (to what extent an individual values some value more than other values. The scores corresponding to each value are then averaged to obtain the per value scores.
A.6 Additional details on The Donation Task
The donation task was created using a template into which different fictional characters from the J.R.R. Tolkien’s universe were inserted. This task was inspired by a charity donation task in ? (?). Figure 12 shows an example of a query for an elf called Círdan, following a topic of “grammar”. A total of 100 different queries are constructed with five races (elves, dwarves, orcs, humans, and hobbits) each with 20 names (10 male and 10 female). Standard names for each race were chosen by taking random names from a list of all characters organized by race. We aimed to avoid specific characters for the beggars (e.g. we imply some unspecified hobbit called Jessamine). However, this could not be entirely achieved due to many names being unique.
We compute the average donated amount and estimate Rank-Order stability separately for each race separately, and then compute the final Rank-Order stability as the average over races. We treat donations to different races similarly to different values in PVQ. This choice was made because one could expect evil characters to donate more to orcs compared to other races, and vice versa. Hence, averaging over all characters might hide relevant trends in donation.
Appendix B Additional experiments and analyses
B.1 Visualization of Mixtral-Instruct value expression
In this section, we visualize the PVQ values expressed by the Mixtral-Instruct model along different contexts and seeds. We use PCA (?, ?) to visualize 250 dimensions (5 seeds x 5 contexts x 10 values) dimensions as two PCA components with $R^{2}=0.29$ and $R^{2}=0.12$ explained variance ratios. We used GPT-4 to classify characters into positive, neutral (more complex), and negative using the following prompt:
     Classify the following characters into positive/neutral/negative.    Create a two column table with the first column being the name of  the character and the second being the classification.    Gandalf  ...  ...  Gimli
Fig. 8 shows the representation of 60 fictional characters classified into positive, neutral (more complex) and negative characters. We can see that positive characters are grouped on the left, negative characters on the right, and neutral in the middle. This shows the Mixtral-Insturct model expressed values in a semantically plausible way.
Figure 8: PCA representation of different fictional characters simulated by the Mixtral-Instruct model. Positive characters (green) are grouped on the left side, negative characters (red) on the right side, and neutral (more complex) characters are in the middle.
B.2 Do simulated personas’ value profiles approach a neutral value profile with longer conversations?
In section 3.4 in the main text, we studied Mixtral-Instruct that was instructed to simulate fictional characters. We observed that Rank-order stability diminished and Ipsative stability stayed the same with longer conversations. This implied that simulated personas’ value profiles moved away from the instructed personas towards some neutral value profile. Here, we experimentally confirm this hypothesis by estimating the distance of simulated personas’ value profiles to an estimated neutral value profile. The distance is computed as Ipsative stability: correlation between the order of values in a simulated individual to those of the neutral value profile.
The neutral profile is estimated as follows. We evaluate the Mixtral-Instruct model without the persona setting instructions and without simulating a conversation (i.e. the questionnaire queries are given straight away). We repeat this process with 50 permutations in the order of suggested answers. To estimate the neutral profile, we average the value ranks over those permutations as shown in the following pseudocode:
     # value_score.shape == (50, 10)  value_scores = evaluate_model("Mixtral-Instruct", n_perm=50)  all_ranks = []  for i in range(50):      # permutation_value_profile.shape == (10)      permutation_scores = value_scores[i]      permuataion_value_ranks = compute_ranks(permuatation_scores)      all_ranks.append(permutation_values_ranks)    # all_ranks.shape == (50, 10)  # neutral_value_profile.shape == (10)  neutral_value_profile = all_rank.mean(axis=0)
Fig. 9 shows the similarity of simulated personas’ value profiles to the default profile, with the Rank-Order stability of simulated individuals between contexts (same as figure 3(a)). We can see that, as Rank-Order stability diminishes, simulated personas’ value profiles move closer to the default profile. This confirms our hypothesis that the diminishing Rank-Order stability is due to the model gradually “ignoring” the persona inducing instruction and moving all simulated personas’ value profiles closer to the neutral one.
Figure 9: Similarity of Mixtral-Instruct simulated fictional characters’ value profiles with the neutral value profile (blue) compared to the Rank-order stability. As conversations gets longer, simulated value profiles move away from the instructed persona toward a neutral one, resulting in lower Rank-order stability.
B.3 Is value expression correlated with behavior on a downstream task?
In section 3.5 in the main text, we studied if models that exhibit more stable value profiles also exhibit more stable behavior on a downstream task. Here, we study if value expression correlates with that behavior. We hypothesize that, for more stable models, simulated personas that exhibited higher universalism and benevolence will also donate more coins. Similarly, simulated personas that exhibited higher power and achievement should donate less.
We compute the correlations between the order of simulated participants in terms of expression of some value (e.g. Universalism) and the donation to each of the four fictional races (a total of 4 correlations), and compute the mean of those correlations. In doing so, the contexts are paired (e.g. the expression of Benevolence following a conversation about grammar is correlated with the amount donated to elves following a conversation about grammar).
Fig. 14 shows the correlation between rank-order of value expression on PVQ and the donation amount on a downstream task. As hypothesized, we can see that for most stable models: Mistral-Instruct-v1, Mixtral-Instruct (both versions), and Qwen-72B donations are correlated with Universalism and Benevolence, and negatively correlated with Power and Achievement. We again observe a trend in model families, with Mixtral, Mistral, and Qwen being more stable than LLaMa-2 and Phi. Having said that, neither model exhibited high correlation ($<0.3$). This experiment implies that, while expected positive and negative correlations between value expression and donation are observed, there is much room for improvement.
B.4 Do LLaMa-2 models exhibit low stability because of the persona induction method?
In the main text, LLaMa-2 chat models exhibited very low stability, but those models are also the only ones (apart from zephyr-7b-beta) which used the prompt template with the system message input as discussed in appendix A and shown in figure 10(b). Furthermore, Mistral-Instruct and Mixtral-Instruct, which showed high stability, used the template without the system message input (Fig. 10(c)). Therefore, we found it relevant to check that LLaMa-2 low stability is not caused by the prompting template but by the model itself.
Figure 13 shows the three LLaMa-2 chat tuned models with the two prompting templates. It compares inducing the persona through the system message (denoted by "*_sys"), as was done in the main text, to inducing it through the user message (denoted by "*_no_sys"), as was done for other models. We can see that neither prompt template enables the LLaMa-2 chat models to exhibit higher stability. This implies that the exhibited low stability is due to the models themselves, and not merely due to the choice of a prompting template.
Figure 10: Rank-order value stability ($Mean\pm SI(\alpha=0.05)$) of chat-tuned LLaMa-2 models when the persona is induces through the system message (as was done in the main text) compared to the user message input (as was done for other models). LLaMa-2 models do not exhibit value stability in either setting. This implies that the low stability is due to the LLaMa models themselves, and not due to the choice of a prompting template.
((a)) Base models.
((b)) Instruction or Chat tuned model with the system input.
((c)) Instruction or Chat tuned model without the system input.
Figure 11: Prompt examples of administering a questionnaire to different models. For tuned models, <system>, <user>, and <assistant> are replaced with specific keywords defined by their fine-tuning. A persona (blue) and a conversation topic (green) are induced. A conversation is simulated (orange). A query from a questionnaire or a downstream task (purple) is given followed by the “query string’ (purple). The query string makes the next token distribution much more skewed towards capital letters denoting an answer. The model generates a distribution for the next token, and the answer is taken as the most probable token from a set of capital letters from A to F.
Figure 12: An example of querying the model in the downstream donation task following a conversation about grammar. The example is the same as that in figure 10(b) with a different query (shown in red).
Table 1: A list of personas in the simulated populations
Population
Personas
Fictional characters
Gandalf, Gollum, Sméagol, Aragorn, Sauron, Saruman, Celeborn, Galadriel, Tom Bombadil, Elrond, Frodo Baggins, Maedhros, Finrod Felagund, Glorfindel, Goldberry, Bilbo Baggins, Smaug, Morgoth, Faramir, Éowyn, Samwise Gamgee, Fëanor, Théoden, Boromir, Túrin Turambar, Thranduil, Beorn, Arwen, Halbarad, Isildur, Gothmog (Balrog), Lungorthin, Celebrimbor, Gil-galad, Meriadoc Brandybuck, Treebeard, Shelob, Radagast, Elendil, Denethor, Éomer, Legolas, Húrin, Thorin Oakenshield, Peregrin Took, Thingol, Gríma Wormtongue, Eärendil, Elwing, Lúthien, Beren, Tuor, Idril, Finwë, Míriel, Ungoliant, Thuringwethil, Melian, Durin’s Bane, Gimli
Real-world persona
Marilyn Monroe, Abraham Lincoln, Nelson Mandela, Queen Elizabeth II, John F. Kennedy, Martin Luther King, Winston Churchill, Donald Trump, Bill Gates, Muhammad Ali, Mahatma Gandhi, Mother Teresa, Christopher Columbus, Charles Darwin, Elvis Presley, Albert Einstein, Paul McCartney, Queen Victoria, Pope Francis, Jawaharlal Nehru, Leonardo da Vinci, Vincent Van Gogh, Franklin D. Roosevelt, Pope John Paul II, Thomas Edison, Rosa Parks, Lyndon Johnson, Ludwig Beethoven, Oprah Winfrey, Indira Gandhi, Eva Peron, Benazir Bhutto, George Orwell, Vladimir Putin, Dalai Lama, Walt Disney, Neil Armstrong, Peter Sellers, Barack Obama, Malcolm X, J.K.Rowling, Richard Branson, Pele, Angelina Jolie, Jesse Owens, John Lennon, Henry Ford, Haile Selassie, Joseph Stalin, Lord Baden Powell
Figure 13: Rank-order value stability ($Mean\pm SI(\alpha=0.05)$) of chat-tuned LLaMa-2 models when the persona is induces through the system message (as was done in the main text) compared to the user message input (as was done for other models). LLaMa-2 models do not exhibit value stability in either setting. This implies that the low stability is due to the LLaMa models themselves, and not due to the choice of a prompting template.
((a)) Universalism
((b)) Benevolence
((c)) Power
((d)) Achievement
Figure 14: Rank-order stability between value expression (on the PVQ questionnaire) and the donation amount (on a downstream task). For more stable modes, donations are correlated with Universalism and Benevolence and negatively correlated with power and achievement.